# Generation of Indoor Open Street Maps for Robot Navigation from CAD Files

### v1

@@ Abstract
The deployment of autonomous mobile robots is predicated on the availability of environmental maps, yet conventional generation via SLAM (Simultaneous Localization and Mapping) suffers from significant limitations in time, labor, and robustness, particularly in dynamic, large-scale indoor environments where map obsolescence can lead to critical localization failures. To address these challenges, this paper presents a complete and automated system for converting architectural Computer-Aided Design (CAD) files into a hierarchical topometric OpenStreetMap (OSM) representation, tailored for robust life-long robot navigation. Our core methodology involves a multi-stage pipeline that first isolates key structural layers from the raw CAD data and then employs an AreaGraph-based topological segmentation to partition the building layout into a hierarchical graph of navigable spaces. This process yields a comprehensive and semantically rich map, further enhanced by automatically associating textual labels from the CAD source and cohesively merging multiple building floors into a unified, topologically-correct model. By leveraging the permanent structural information inherent in CAD files, our system circumvents the inefficiencies and fragility of SLAM, offering a practical and scalable solution for deploying robots in complex indoor spaces. The software is encapsulated within an intuitive Graphical User Interface (GUI) to facilitate practical use. The code and dataset are available at https://github.com/jiajiezhang7/osmAG-from-cad.
@@ I Introduction
The proliferation of autonomous mobile robots in indoor environments, such as hospitals and warehouses, demands robust and persistent navigation capabilities, for which an accurate environmental map is a fundamental prerequisite. For practical, long-term deployments—often termed ”life-long” navigation—maintaining a correct map within a dynamic environment is paramount. Traditional on-site mapping methods, predominantly based on Simultaneous Localization and Mapping (SLAM) [1], exhibit significant drawbacks. The map-building process is often labor-intensive and time-consuming [2, 3]. Moreover, the resulting maps, especially dense metric formats like 2D occupancy grids [4] or 3D point clouds [5], are subject to environmental changes (e.g., rearranged furniture), which can lead to catastrophic localization failures. Their substantial data volume also imposes high storage and computational overhead, rendering them inefficient for long-horizon planning in large-scale settings [6, 7]. Moreover, during data collection for SLAM, often not all rooms are accessible, resulting in incomplete maps.
@@
To overcome these challenges, we shift the paradigm from on-the-fly mapping to leveraging existing architectural information. This paper presents a mostly automated pipeline that converts standard 2D Computer-Aided Design (CAD) files into an enhanced OpenStreetMap (OSM) [10] representation. The generated map is not merely a geometric blueprint but is enriched with topometric details and a hierarchical structure, encoding both the metric layout and semantic relationships between spaces (e.g., rooms, corridors). This enhanced OSM serves as a robust and persistent foundation for advanced robotic tasks, directly enabling centimeter-level localization [11] and efficient, semantically-aware global path planning [12] and navigation, and thus is a core part of mobile manipulation and embodied AI.
@@ III-A CAD Preprocessing and Occupancy Grid Generation
The initial stage of our pipeline converts the raw architectural CAD floor plan into a standardized binary occupancy grid, which serves as the foundation for all subsequent analyses. This conversion involves two primary steps. First, we perform geometry extraction by filtering layers to isolate permanent structural features essential for navigation, such as walls, columns, and windows. To automate this process, we employ a keyword-based filtering mechanism that targets layer names conforming to established architectural standards (e.g., ”A-WALL,” ”A-STAIR”), such as the US National CAD Standard (NCS) [26] and ISO 13567 [27]. For files that deviate from these conventions, minimal manual layer selection is required prior to processing.
Naively, one would attempt to then use the vector graphics information in the CAD to directly extract the area ploygons for osmAG. But often, rooms and areas in CAD are not watertight, i.e. the lines for a room drawn by the architect do not always from closed polygons. Thus, in the second step, the filtered vector geometry is rasterized into a high-resolution binary image at a precise metric scale. During rasterization, we apply a morphological processing algorithm to thicken wall structures and bridge minor discontinuities. This enhancement ensures the floor plan’s free space is partitioned into well-defined, enclosed connected components, producing a clean occupancy grid. This grid, with unambiguously delineated structures, is a vital prerequisite for robust topological segmentation by the subsequent AreaGraph algorithm.
@@ III-B Topological Graph Generation and OSM Transformation
4:$V\leftarrow$ ComputeVoronoiDiagram($G$)
5:$T\leftarrow$ ExtractTopologicalSkeleton($V$)
6:$R\leftarrow$ AlphaShapeSegmentation($T$, $\alpha$)
7:$AG\leftarrow$ BuildAreaGraph($R$)
@@
9:$AG.rooms\leftarrow$ RemoveDuplicatePolygons($AG.rooms$)
10:$AG.rooms\leftarrow$ MergeSmallRooms($AG.rooms$, $A_{min}$, $d_{max}$)
11:$P_{preserve}\leftarrow$ CollectPassageEndpoints($AG.passages$)
@@
17:$nodes\leftarrow\emptyset$, $ways\leftarrow\emptyset$
18:$id_{counter}\leftarrow-1$
@@ III-C Semantic Text-to-Room Association
Figure 5: Results of Hierarchical Multi-Floor Map Fusion. Figures 5(a),5(b),5(c) depict three distinct hierarchical levels of a ShanghaiTech building as rendered in the standard OpenIndoor viewer[9]. Figure 5(d) showcases an enhanced OSM map visualized in Rviz after being processed by osmAG parser [12]. Within this visualization, the red line segments represent vertical passages (e.g. elevators and stairs) that topologically connect adjacent floors. This enhanced representation is specifically designed to enable cross-floor mobile robot localization [11] and path planning [12], and its effectiveness for both applications has been validated, as demonstrated in the accompanying video.
@@ III-D Hierarchical Multi-Floor Map Fusion
The culminating step of the topometric processing pipeline involves serializing the refined topological graph into the OpenStreetMap (OSM) XML format, a process distinguished by its precise geo-referencing capabilities. By defining the WGS84 coordinates (longitude and latitude) for a designated origin point alongside the map’s metric resolution, the system generates an OSM file where all geometric elements are situated within a global reference frame. This transformation yields a spatially accurate, interoperable, and permanent map artifact that can be seamlessly integrated with the global OSM dataset. Within this structure, each navigable space is encoded as a closed OSM way tagged as indoor=room. Passages connecting these spaces are likewise represented as way elements, but are assigned the custom tag osmAG:type=passage and include attributes identifying the two areas they link. This tagging scheme ensures full compatibility with the official OSM standard while embedding the enhanced topological information essential for advanced mobile robot navigation.
@@ IV EXPERIMENTS
To validate the effectiveness, robustness, and scalability of our proposed automated map generation system, we conducted a series of experiments on a diverse collection of architectural CAD floor plans. As our work presents a novel end-to-end pipeline for converting CAD files into hierarchical topometric OSM maps, there are no existing methods for direct performance comparison. Therefore, our evaluation focuses on quantifying the accuracy and efficiency of our system against manually annotated ground truth.
@@ V Conclusion and Discussion
The efficacy and scalability of our system were validated through the successful generation of a map for a large-scale ($9,000~{}m^{2}$) building and across a diverse set of real-world CAD files, achieving high accuracy in both topological segmentation and semantic labeling. By shifting the paradigm from on-site sensing to utilizing pre-existing architectural data, our work provides a practical and efficient solution that enables robust, life-long robot navigation, precise localization, and semantically-aware path planning, as also shown in the video accompanying this paper.
Future work will focus on enhancing the system’s autonomy by developing an adaptive parameter selection mechanism for segmentation and extending the algorithm to natively handle layouts with disjointed navigable spaces. Moreover, the foundational principles of our pipeline are generalizable and could be adapted to construct osmAG maps from other modalities, such as large-scale point clouds. These advancements, coupled with potential integration with richer data sources like BIM, will further advance the creation of persistent, life-long maps for robotics.

### v2

@@ Abstract
The deployment of autonomous mobile robots relies on environmental maps, which are conventionally generated via SLAM (Simultaneous Localization and Mapping). While effective for initial mapping, these maps face a persistence challenge in long-term applications, as they can become obsolete in dynamic indoor environments, leading to critical localization failures. To address this, this paper presents a complete and automated system for converting architectural Computer-Aided Design (CAD) files into a hierarchical topometric OpenStreetMap (OSM) representation, tailored for robust life-long robot navigation. Our core methodology involves a multi-stage pipeline that first isolates key structural layers from the raw CAD data and then employs an AreaGraph-based topological segmentation to partition the building layout. This process yields a semantically rich map, further enhanced by automatically associating textual labels and cohesively merging multiple floors into a unified model. By leveraging the permanent structural information inherent in CAD files, our system provides a stable and persistent map foundation, offering a practical and scalable solution to the map maintenance challenge in long-term robot deployments. The software is encapsulated within an intuitive Graphical User Interface (GUI) to facilitate practical use. The code and dataset are available at https://github.com/jiajiezhang7/osmAG-from-cad.
@@ I Introduction
The proliferation of autonomous mobile robots in indoor environments, such as hospitals and warehouses, demands robust and persistent navigation capabilities, for which an accurate environmental map is a fundamental prerequisite. For practical, long-term deployments—often termed ”life-long” navigation—the challenge extends beyond initial map creation to the critical task of map maintenance within dynamic environments. Traditional on-site mapping methods, predominantly based on Simultaneous Localization and Mapping (SLAM) [1], exhibit significant drawbacks for persistent autonomy. The map-building process is not only labor-intensive and time-consuming [2, 3], but the resultant maps, especially dense metric formats like 2D occupancy grids [4] or 3D point clouds [5], are subject to environmental changes (e.g., rearranged furniture), which can lead to critical localization failures.
@@
To overcome these challenges, we shift the paradigm from on-the-fly, sensor-driven mapping to leveraging the permanent structural information inherent in architectural designs. We posit that Computer-Aided Design (CAD) files provide the authoritative ground truth of a building’s layout, which, unlike sensor-based maps, is invariant to ephemeral changes. This paper presents a mostly automated pipeline that converts standard 2D CAD files into an enhanced OpenStreetMap (OSM) [8] representation. The generated map is not merely a geometric blueprint but is enriched with topometric details and a hierarchical structure, serving as a robust and persistent foundation for advanced robotic tasks.
@@ III-A CAD Preprocessing and Occupancy Grid Generation
The initial stage of our pipeline converts the raw architectural CAD floor plan into a standardized binary occupancy grid, which serves as the foundation for all subsequent analyses. This conversion involves several steps, as depicted in Fig. 2. First, we perform geometry extraction by filtering layers to isolate permanent structural features essential for navigation, such as walls, columns, and windows. To automate this process, we employ a keyword-based filtering mechanism that targets layer names conforming to established architectural standards (e.g., ”A-WALL,” ”A-STAIR”), such as the US National CAD Standard (NCS) [24] and ISO 13567 [25]. For files that deviate from these conventions, minimal manual layer selection is required.
Following layer filtering, the system performs a Geo-referenced Vectorization step. Here, the clean vector graphics are exported as an SVG file, and critically, the transformation parameters relating the vector coordinates to a real-world metric scale are preserved. This process also computes the Image Bounds, which define the precise physical dimensions and origin of the floor plan. In the final step, this geo-referenced vector data is rasterized into a high-resolution binary image at the correct metric scale. During rasterization, we apply a morphological processing algorithm to thicken wall structures and bridge minor discontinuities, ensuring the floor plan’s free space is partitioned into well-defined, enclosed connected components. This produces a clean occupancy grid, a vital prerequisite for robust topological segmentation.
@@ III-B Topological Graph Generation and OSM Transformation
[…]
@@
[…]
@@
[…]
@@ III-C Semantic Text-to-Room Association
Figure 5: Results of Hierarchical Multi-Floor Map Fusion. Figures 5(a),5(b),5(c) depict three distinct hierarchical levels of a ShanghaiTech university building as rendered in the standard OpenIndoor viewer[7]. Figure 5(d) showcases an enhanced OSM map visualized in Rviz after being processed by osmAG parser [9]. Within this visualization, the red line segments represent vertical passages (e.g. elevators and stairs) that topologically connect adjacent floors. This enhanced representation is specifically designed to enable cross-floor mobile robot localization [18] and path planning [9], and its effectiveness for both applications has been validated, as demonstrated in the accompanying video.
@@ IV EXPERIMENTS
To validate the accuracy and efficiency of our proposed automated map generation system, we conducted a series of experiments on a diverse collection of architectural CAD floor plans. As our work presents a novel end-to-end pipeline for converting CAD files into hierarchical topometric OSM maps, a direct comparison to other map generation methods is not feasible. Therefore, our evaluation focuses on quantifying the performance of the core components of our system—namely the topological segmentation and semantic association—against manually annotated ground truth. The primary goal is to validate the fidelity of the generated map artifact itself. The effectiveness of the resulting osmAG map structure for downstream robotic tasks has been validated in prior work, which demonstrated its utility for robust, centimeter-level localization [18] and efficient, semantically-aware global path planning [9].
@@ V Conclusion and Discussion
The efficacy and scalability of our system were validated through the successful generation of a map for a large-scale ($9,000~m^{2}$) building and across a diverse set of real-world CAD files, achieving high accuracy in both topological segmentation and semantic labeling. By shifting the paradigm from on-site sensing to utilizing pre-existing architectural data, our work provides a practical and efficient solution for creating persistent map backbones that enable robust life-long robot navigation, precise localization [18], and semantically-aware path planning [9].
While our system demonstrates a robust pipeline, we acknowledge limitations that offer avenues for future work. The current reliance on keyword-based filtering for layer selection requires manual intervention for non-compliant drawings; this could be overcome by exploring machine learning for automatic layer classification. Similarly, the heuristically set parameters for segmentation could be replaced by an adaptive selection mechanism to enhance generalizability. A more fundamental challenge is that architectural plans can become outdated. We propose that our generated map can serve as a high-fidelity static baseline, which can be augmented with dynamic layers from on-site sensors. This hybrid approach, merging the permanence of architectural data with real-time awareness, represents a promising direction for creating truly persistent, life-long maps for robotics.